OpenStudy (anonymous):

I need to find the angle... a mini coaster moves 300 m horizontally and then rises 200 m at an angle of 20° above the horizontal then 200 m at an angle of 34° downward (below the horizontal). What is the angle counterclockwise from the +x-axis? (Assume the initial horizontal motion of the car to be along the +x axis.)

OpenStudy (agent0smith):

Draw a diagram first. Then you'll have to find the horiz and vert components of each vector, and use that to get the resultant vector (and then the angle).

OpenStudy (anonymous):

I did. But I am getting the wrong answer. Can you show me how to get the angle. I don't need the answer. I just want to know how to get it .

OpenStudy (agent0smith):

Well, your net horizontal vector is = 300 + 200cos20 + 200cos34 net vertical = 200sin20 - 200 sin34 now use tangent = vert/horizontal (you'll have to subtract the angle from 360 since you want the counterclockwise angle)

OpenStudy (loser66):

@agent0smith can you please draw a diagram?

OpenStudy (anonymous):

The computer said that my answer is wrong

OpenStudy (anonymous):

I did same thing...I am not getting the angle

OpenStudy (agent0smith):

Show some work so we can see what you're doing wrong. Angle should be around 330 degrees or so (just from looking at the pic, remember you're measuring clockwise from the positive x axis) 

OpenStudy (anonymous):

it says counter clockwise

OpenStudy (agent0smith):

Yeah, that's what i meant.

OpenStudy (anonymous):

I understand how to do it, but the problem that I pm you... I am getting the same answer...358 degrees...